Differential this vibration equation from dynamic vibration absorber (dva) system. if x2 is the output of this system.
Upper body bos Lower body 777777 Base
Answer #1

3 1. 1 h 0 → Use Nodal Methodi (a) No. of nodes = No. of displacements = 3 (x,,M2,2) (b) Take an additional node which is - aat alt > me de - m, d + ki o bun-ki center +6,4,-CZ + K₂ dl - C + C₂1,- Calz=0 → mi de mi ay +(x,*Ka) de K dz + (( +(2)% -6,272 t dez Add equation 0 and equation . mi se more + (xv+Kia) +((i+C2)2, -C, 2-%d722 -C2% + m₂ $242 - m2 2 2 2 2 + K₂ d 2 = 2